Show that the series n=1∑∞ x/(1+n2x2) is uniformly convergent in [α,1] for any α>1.


1
Expert's answer
2021-06-15T16:16:13-0400

"\sum_{n=1}^{\infty} \frac{x}{1+n^2x^2}
,x\in [\alpha, 1] \text{for any}\alpha<
1.\\
Then,\\
f_n=\frac{x}{1+n^2x^2}
\to0\space as\space n\to \infty\\
M_n=sup_{x\in[\alpha, 1]}|f_n-f|\\
=sup_{x\in[\alpha, 1]}|\frac{x}{1+n^2x^2}-0|\\
=sup_{x\in[\alpha, 1]}|\frac{x}{1+n^2x^2}|\\
\text{To get this value, we find the value of x where the given sequence of function is maximum.}\\
f'_n=\frac{1-n^2x^2}{(1+n^2x^2)^2}\\
Now,\\
f'_n=0\implies x=\frac{1}{n}\\
f''_n=\frac{2n^2x(n^2x^2-3)}{(1+n^2x^2)^3}\\
f''_n(\frac{1}{n})<0\\
\implies at \space x=\frac{1}{n}, \space
M_n \space
\text{attain maximum}.\\
M_n=sup_{x\in[\alpha, 1]}|\frac{x}{1+n^2x^2}|\\
=|\frac{\frac{1}{n}
}{1+n^2\frac{1}{n^2}}|\\
=\frac{1}{2n} \to 0\space as \space
n\to \infty\\
\text{ Therefore, by Weierstrass M-test the given sequence of function converge. }"